OpenStudy (anonymous):

lim (sinx/x) when x-->0 means x goes to zero when we substitute zero, we get 0/0 so we have to use hopital's rule, that's mean diffrentiate, so we get (x cosx-sinx)/x^2 when we substitute zero, we get 0/0 again the book says it equals one !! do you have any explanation did I make a mistake in any step? thanx

myininaya (myininaya):

hey lion you do derivative of top/derivative of bottom

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{\sin(x)}{x}=\lim_{x \rightarrow 0}\frac{(\sin(x))'}{(x)'}\]

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{\cos(x)}{1}\]

OpenStudy (sburchette):

You don't differentiate the same way you would when using L'Hopital's rule. You used the quotient rule which is correct in most cases. However, when using L'Hopital's rule, you differentiate the top and bottom independently. So you will get\[\lim_{x \rightarrow 0} [(\sin x)/x]\]Now you differentiate sinx and x.\[\lim_{x \rightarrow 0} \cos x/1\]Substituting the limit you get\[\cos (0)=1\]

OpenStudy (anonymous):

That awesome work guys thank you so much I got the idea completely